OpenStudy (anonymous):

If x+3/3=y+2/2 , then x/3 = ____. Question 17 options: a) y + 1 b) y/3 c) y/2 d) y – 1

OpenStudy (rsadhvika):

\(\large \frac{x+3}{3} = \frac{y+2}{2}\) \(\large \frac{x}{3}+1 = \frac{y+2}{2}\) \(\large \frac{x}{3} = \frac{y+2}{2}-1\)

OpenStudy (rsadhvika):

simplify

OpenStudy (anonymous):

what am i simpifying?